During 2016, Seth Britain Lab supplied hospitals with a comprehensive diagnostic kit for $125. At a volume of 90,000 kits, Seth had fixed costs of $1,140,000 and a profit before income taxes of $300,000. Due to an adverse legal decision, Seth's 2017 liability insurance increased by $1,440,000 over 2016. Assuming the volume and other costs are unchanged, what should the 2017 price be if Seth is to make the same $300,000 profit before income taxes? (CPA adapted)

$128.30.

$141.00.

$160.30.

$250.00
Answer #1
Sales in 2016 11250000 =90000*125
Total variable cost in 2016 9810000 =11250000-1140000-300000
Unit variable cost in 2016 109 =9810000/90000
Required sales 12690000 =(90000*109)+1140000+1440000+300000
2017 selling price 141.00 =12690000/90000
Option B $141.00 is correct